# Benchmarking Algorithms for Federated Domain Generalization

### v1

@@ Abstract
While prior domain generalization (DG) benchmarks consider train-test dataset heterogeneity, we evaluate Federated DG which introduces federated learning (FL) specific challenges. Additionally, we explore domain-based heterogeneity in clients’ local datasets—a realistic Federated DG scenario. Prior Federated DG evaluations are limited in terms of the number or heterogeneity of clients and dataset diversity. To address this gap, we propose an Federated DG benchmark methodology that enables control of the number and heterogeneity of clients and provides metrics for dataset difficulty. We then apply our methodology to evaluate 13 Federated DG methods, which include centralized DG methods adapted to the FL context, FL methods that handle client heterogeneity, and methods designed specifically for Federated DG. Our results suggest that despite some progress, there remain significant performance gaps in Federated DG particularly when evaluating with a large number of clients, high client heterogeneity, or more realistic datasets. Please check our extendable benchmark code here: https://github.com/inouye-lab/FedDG_Benchmark.
@@ 1 Introduction
Domain generalization (DG) (Blanchard et al., 2011) formalizes a special case of train-test heterogeneity in which the training algorithm has access to data from multiple source domains but the ultimate goal is to perform well on data from an unseen test domain—i.e., a type of out-of-distribution generalization instead of the standard in-distribution generalization. While most prior DG work focuses on centralized algorithms, another natural context is federated learning (FL) (Konečnỳ et al., 2016), which is a distributed machine learning context that assumes each client or device owns a local dataset. These local datasets could exhibit heterogeneity, which we call client heterogeneity (e.g., class imbalance between clients). Although train-test heterogeneity (in DG) and client heterogeneity (in FL) are independent concepts, both could be naturally defined in terms of domain datasets. For example, suppose a network of hospitals aimed to use FL to train a model to predict a disease from medical images. Because the equipment is different across hospitals, it is natural to assume that each hospital contains data from different domains or environments (or possibly a mixture of domains if it is a large hospital)—this is a case of domain-based client heterogeneity. Yet, the trained model should be robust to changes in equipment within a hospital or to deployment in a new hospital that joins the network—both are cases of domain-based train-test heterogeneity. The interaction between these types of heterogeneity produces new algorithmic and theoretic challenges yet it may also produce new insights and capabilities. Solutions to Federated DG with both types of heterogeneity could increase the robustness and usefulness of FL approaches because the assumptions more naturally align with real-world scenarios rather than assuming the datasets are i.i.d. This could enable training on partial datasets, increase robustness of models to benign spatial and temporal shifts, and reduce the need for retraining.
@@
In the centralized regime, various approaches have been proposed for DG, including feature selection, feature augmentation, etc. Most of these methods are not applicable in the FL regime which poses unique challenges. In the FL regime, client heterogeneity has long been considered a statistical challenge since FedAvg (McMahan et al., 2017), where it experimentally shows that FedAVG effectively mitigate some client heterogeneity. There are many other extensions based on the FedAvg framework tackling the heterogeneity among clients in FL, for example using variance reduction method (Karimireddy et al., 2020). An alternative setup in FL, known as the personalized setting, aims to learn personalized models for different clients to tackle heterogeneity, for example Hanzely et al. (2020). However, none of these works consider model robustness under domain shift between training and testing data. Recently, a few works in the FL regime tackling DG (Liu et al., 2021; Zhang et al., 2021; Nguyen et al., 2022; Tenison et al., 2022) have been proposed, however their evaluations are limited in the following senses: 1) The evaluation datasets are limited in the number and diversity of domains. 2) The evaluations are restricted to the case when the number of clients is equal to the number of domains, which may be an unrealistic assumption (e.g., a hospital that has multiple imaging centers or a device that is used in multiple locations). The case when clients number might be massive are of both theoretical and application interests. 3) None of the works consider the influence of the effect of the number of communication rounds. We provide an overview of the tasks inTable 1, considering both the heterogeneity between training and testing datasets (standard vs. domain generalization) and among clients (domain client heterogeneity). While some studies have addressed the standard supervised learning task, there is a need for a fair evaluation to understand the behavior of domain generalization algorithms in the FL context under those new challenges.
@@
There are several benchmark datasets available for evaluating domain generalization (DG) methods in the centralized setting. These benchmarks, such as DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021), provide multiple datasets that are suitable for assessing the performance of DG algorithms. However, they did not explicitly consider the unique challenges that arise in the federated learning (FL) setting. On the other hand, there are also benchmarks specifically designed for FL. For instance, the LEAF benchmark (Caldas et al., 2018) provides a standardized framework for evaluating FL algorithms. It includes several datasets from various domains and allows researchers to assess the performance of their algorithms in a realistic FL scenario. Another benchmark for FL is PFLBench (Chen et al., 2022), which focuses on evaluating personalized FL methods. PFLBench provides 12 datasets containing various applications. Though these FL-based benchmarks consider statistical heterogeneity, they fail to consider the DG task adequately. Moreover, the level of statistical heterogeneity present in these datasets is insufficient for proper DG evaluation. In summary, DG benchmarks do not consider FL challenges, and FL benchmarks do not consider DG challenges.
@@
Major contributions: We develop a benchmark methodology for evaluating Federated DG with various client heterogeneity contexts and diverse datasets, and we evaluate representative Federated DG approaches with this methodology. 1) We propose the first Federated DG benchmark methodology including four important dimensions of the experimental setting (see Section 3). 2) We propose a standardized definition of domain-based client heterogeneity that is unique to the FL context and interpolates between domain homogeneity and domain separation (see Fig. 1) while controlling class imbalance. In particular, we develop a novel method to split any dataset with domain labels across any number of clients (see subsection 3.2). 3) We compare three broad approaches to Federated DG: centralized DG methods naïvely adapted to FL setting, FL methods developed for client heterogeneity (e.g., class imbalance), and recent methods specifically designed for Federated DG. Our results indicate that there still exist significant gaps and open research directions in Federated DG. 4) We will release an extendable open-source library for evaluating Federated DG methods.
Notation
@@
2 Approaches to Federated Domain Generalization
We first briefly review the DG problem, extend to the Federated DG problem, and explain domain-based client heterogeneity. Then, we discuss three categories of Federated DG approaches.
@@ 2.1 Problem Background and Setup
Domain generalization (train-test heterogeneity)
Unlike standard ML which assumes the train and test are independent and identically distributed (i.i.d.), the ultimate goal of DG is to minimize the average or worst-case loss of the test domain distributions when only samples from the train domain distributions are given. Formally, given a set of train domain distributions $\{p_{d}:d\in\mathcal{D}_{\text{train}}\}$, minimize the average or worst case loss over test domain distributions, i.e.,
@@
where $\mathcal{L}(\theta;p_{d})=\mathbb{E}_{(x,y)\sim p_{d}}[\ell(x,y;\theta)]$ where $\ell$ is a per-sample loss function such as squared or cross-entropy loss. The key challenge in DG is that the train and test domain distributions are disjoint, i.e., $\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{test}}=\emptyset$, and thus, the method must be able to generalize beyond the train domain distributions to perform well on the test domain distributions. The naïve approach is to simply ignore the domains and perform empirical risk minimization (ERM) on all the training data—which is actually a challenging baseline to outperform in practice (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021).
Federated DG
Federated DG adds a layer of complexity because now the domain distribution samples are not centrally located. Instead, each FL client can update their local model based only on their local client distribution $p_{c}$ and then pass their local parameters $\theta_{c}$ to a central server, which will aggregate the local models and broadcast the model to all clients. The FL problem can be abstracted as follows:
@@
$\forall c\in[C],\,\,\underbrace{\theta_{c}=\texttt{Optimize}(\mathcal{L}(\theta;p_{c}),\theta_{\text{init}}=\theta_{\text{global}})}_{\text{Locally optimize given local distribution $p_{c}$}}\quad\text{and}\quad\underbrace{\theta_{\text{global}}=\texttt{Aggregate}(\theta_{1},\theta_{2},\cdots,\theta_{C})}_{\text{Aggregate client model parameters on server}}\,,$
@@
where the client distributions may be homogeneous (i.e., $\forall(c,c^{\prime}),p_{c}=p_{c^{\prime}}$) or heterogeneous (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$), Optimize minimizes an objective $\mathcal{L}(\theta;p_{c})$ initialized at $\theta_{\text{init}}$, and Aggregate aggregates the client model parameters, where the most common aggregator is simply a (weighted) average of the client parameters, which corresponds to FedAvg (McMahan et al., 2017).
Domain-based client heterogeneity
@@
While client heterogeneity (i.e., $\exists c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$) is often expressed as label imbalance, i.e., $p_{c}(y)\neq p_{c^{\prime}}(y)$, we make a domain-based client heterogeneity assumption that each client distribution is a (different) mixture of train domain distributions, i.e., $p_{c}(x,y)=\sum_{d\in\mathcal{D}_{\text{train}}}w_{c,d}\,\,p_{d}(x,y)$ where $w_{c,d}$ is the weight of the $d$-th domain for the $c$-th client. At one extreme, FL with i.i.d. data would be equivalent to the mixture proportions being the same across all clients, i.e., $\forall c,c^{\prime},d,w_{c,d}=w_{c^{\prime},d}$, which we call the homogeneous setting. On the other extreme, if the number of clients and number of domains are equal ($D=C$), the domain weights could be disjoint between clients such that each client “owns” one or more domains, i.e., for any domain $d\in[D],$ $w_{c,d}>0\Leftrightarrow w_{c^{\prime},d}=0,\forall c^{\prime}\neq c$, which we call domain separation (in subsection 3.2, we extend the domain separation to the case $D\neq C$). Finally, we use $\lambda$ to denote an interpolation parameter between these two extremes: homogeneous case $\lambda=1$ and domain separation case $\lambda=0$. We give details of an explicit algorithm to use in practice for splitting datasets for evaluation in subsection 3.2 which covers two extremes and interpolation cases $\lambda\in(0,1)$. Table 1 summarizes the train-test heterogeneity in DG and the domain-based client heterogeneity from the FL context, where we focus on Federated DG.
@@
Table 1: Domain-based Federated DG considers the domain heterogeneity both among the clients’ local datasets (rows) and between the train and test datasets (columns). This paper focuses on the DG setting (right column).

### v2

@@ Abstract
While prior federated learning (FL) methods mainly consider client heterogeneity, we focus on the Federated Domain Generalization (DG) task, which introduces train-test heterogeneity in the FL context. Existing evaluations in this field are limited in terms of the scale of the clients and dataset diversity. Thus, we propose a Federated DG benchmark that aim to test the limits of current methods with high client heterogeneity, large numbers of clients, and diverse datasets. Towards this objective, we introduce a novel data partition method that allows us to distribute any domain dataset among few or many clients while controlling client heterogeneity. We then introduce and apply our methodology to evaluate $14$ DG methods, which include centralized DG methods adapted to the FL context, FL methods that handle client heterogeneity, and methods designed specifically for Federated DG on $7$ datasets. Our results suggest that, despite some progress, significant performance gaps remain in Federated DG, especially when evaluating with a large number of clients, high client heterogeneity, or more realistic datasets. Furthermore, our extendable benchmark code will be publicly released to aid in benchmarking future Federated DG approaches.
@@ 1 Introduction
Domain generalization (DG) (Blanchard et al., 2011) formalizes a special case of train-test heterogeneity in which the training algorithm has access to data from multiple source domains but the ultimate goal is to perform well on test data coming from a different distribution from training distribution—i.e., a type of out-of-distribution generalization instead of the standard in-distribution generalization. While most prior DG work focuses on centralized algorithms, another natural context is federated learning (FL) (Konečnỳ et al., 2016), which is a distributed machine learning context that assumes each client or device owns a local dataset. These local datasets could exhibit heterogeneity, which we call client heterogeneity (e.g., class imbalance between clients). Although train-test heterogeneity (in DG) and client heterogeneity (in FL) are independent concepts, both could be naturally defined in terms of domain datasets. For example, suppose a network of hospitals aimed to use FL to train a model to predict a disease from medical images. Because the equipment is different across hospitals, it is natural to assume that each hospital contains data from different domains or environments (or possibly a mixture of domains if it is a large hospital)—this is a case of domain-based client heterogeneity. Yet, the trained model should be robust to changes in equipment within a hospital or to deployment in a new hospital that joins the network—both are cases of domain-based train-test heterogeneity. The interaction between these types of heterogeneity produces new algorithmic and theoretic challenges yet it may also produce new insights and capabilities. Solutions to Federated DG with both types of heterogeneity could increase the robustness and usefulness of FL approaches because the assumptions more naturally align with real-world scenarios rather than assuming the datasets are i.i.d. This could enable training on partial datasets, increase robustness of models to benign spatial and temporal shifts, and reduce the need for retraining.
@@
In the centralized regime, various approaches have been proposed for DG, including feature selection, feature augmentation, etc. Most of these methods are not applicable in the FL regime which poses unique challenges. In the FL regime, client heterogeneity has long been considered a statistical challenge since FedAvg (McMahan et al., 2017), where it experimentally shows that FedAvg effectively mitigate some client heterogeneity. There are many other extensions based on the FedAvg framework tackling the heterogeneity among clients in FL, for example using variance reduction method (Karimireddy et al., 2020). An alternative setup in FL, known as the personalized setting, aims to learn personalized models for different clients to tackle heterogeneity, for example Hanzely et al. (2020). There are also unsupervised FL methods tackling domain translation instead of classification (Zhou et al., 2022; Wang et al., 2023). However, none of these works consider model robustness under domain shift between training and testing data. Recently, a few works in the FL regime tackling DG (Liu et al., 2021; Zhang et al., 2021; Nguyen et al., 2022; Tenison et al., 2022b) have been proposed, however their evaluations are limited in the following senses: 1) The evaluation datasets are limited in the number and diversity of domains. 2) The evaluations are restricted to the case when the number of clients is equal to the number of domains, which may be an unrealistic assumption (e.g., a hospital that has multiple imaging centers or a device that is used in multiple locations). The case when clients number might be massive are of both theoretical and application interests. 3) None of the works consider the influence of the effect of the number of communication rounds. We provide an overview of the tasks in Table 1, considering both the heterogeneity between training and testing datasets (standard vs. domain generalization) and among clients (domain client heterogeneity). While some studies have addressed the standard supervised learning task, there is a need for a fair evaluation to understand the behavior of domain generalization algorithms in the FL context under those new challenges.
@@
There are several benchmark datasets available for evaluating domain generalization (DG) methods in the centralized setting. These benchmarks, such as DomainBed (Gulrajani and Lopez-Paz, 2020) and WILDS (Koh et al., 2021), provide multiple datasets that are suitable for assessing the performance of DG algorithms. However, they did not explicitly consider the unique challenges that arise in the federated learning (FL) setting. On the other hand, there are also benchmarks specifically designed for FL. For instance, the LEAF (Caldas et al., 2018) and FLamby (Terrail et al., 2022) benchmark provides a standardized framework for evaluating FL algorithms. They include several datasets from various domains and allows researchers to assess the performance of their algorithms in a realistic FL scenario. Another benchmark for FL is PFLBench (Chen et al., 2022), which focuses on evaluating personalized FL methods. PFLBench provides 12 datasets containing various applications. Though these FL-based benchmarks consider statistical heterogeneity, they fail to consider the DG task adequately. Moreover, the level of statistical heterogeneity present in these datasets is insufficient for proper DG evaluation. In summary, DG benchmarks do not consider FL challenges, and FL benchmarks do not consider DG challenges.
@@
Major contributions: We develop a benchmark methodology for evaluating Federated DG with various client heterogeneity contexts and diverse datasets, and we evaluate representative Federated DG approaches with this methodology. 1) We develop a novel method to partition dataset across any number of clients that is able to control the heterogeneity among clients. (see Section 3). 2) We propose the first Federated DG benchmark methodology including four important dimensions of the experimental setting (see Section 4). 3) We compare three broad approaches to Federated DG: centralized DG methods naïvely adapted to FL setting, FL methods developed for client heterogeneity (e.g., class imbalance), and recent methods specifically designed for Federated DG on $7$ diverse datasets. Our results indicate that there still exist significant gaps and open research directions in Federated DG. 4) We release an extendable open-source library for evaluating Federated DG methods.
@@
2 Background: Federated Domain Generalization Methods
In this section, we briefly introduce the problem backup and setup. Furthermore, evaluation on FL regimes often requires data partition methods to distribute training data to each client. Thus we also introduce the existing partition methods in this section.
@@ 2.1 Problem Background and Setup
Domain generalization: Unlike standard ML which assumes the training and test data are independent and identically distributed (i.i.d.), the ultimate goal of DG is to minimize the average or worst-case loss of test domain distributions using only samples from the training domain distributions. Formally, given a set of training domain distributions $\{p_{d}:d\in\mathcal{D}_{\text{train}}\}$, minimize the average or worst case loss over test domain distributions, i.e.,
@@
where $\mathcal{L}(\theta;p_{d})=\mathbb{E}_{(x,y)\sim p_{d}}[\ell(x,y;\theta)]$ where $\ell$ is a per-sample loss function such as squared or cross-entropy loss. Those two objectives are the ultimate goal for domain generalization. In the real-world setting, we never have access to $\mathcal{D}_{\text{test}}.$ There are works establishing objectives to approximate those two objectives. For example, Fish (Shi et al., 2021) is constructed using average training loss with a penalty, and IRM (Arjovsky et al., 2019), GroupDRO (Sagawa et al., 2019) are inspired from the worst-case loss over a set of data distributions constructed from the training domains.
There are mainly two kinds of DG test scenarios. The first scenario is the general DG where the test dataset form a new domain. Another scenario is called subpopulation shift where the test set is a subpopulation of the training distributions with the goal of better performance on the worst-case domain [^1] In Wilds (Koh et al., 2021), they treat subpopulation shift as another kind of train-test heterogeneity. DG is challenging where it breaks the i.i.d. assumptions in traditional machine learning.
Federated DG: Federated DG represents an intuitive progression from centralized DG. Distributed learning is pivotal in numerous real-world applications, and its inherent architecture often results in multi-domain data. The quest for improved generalization is instinctive in this context. However, Federated DG introduces add two layers of complexity. First, the heterogeneous client distribution are harder to converge even without considering the train-test heterogeneity (Zhao et al., 2018; McMahan et al., 2017; Karimireddy et al., 2020; Yu et al., 2019; Basu et al., 2019; Wang et al., 2019; Li et al., 2019). Second, privacy considerations and communication constraints typically prohibit the direct comparison of data across domains, thereby limiting the applicability of many conventional centralized DG methods.
Formally, the FL problem can be abstracted as follows:
@@
$\forall c\in[C],\,\,\underbrace{\theta_{c}=\texttt{Opt}(\mathcal{L}(\theta;p_{c}),\theta_{\text{init}}=\theta_{\text{global}})}_{\text{Locally optimize given local distribution $p_{c}$}}\;\text{and}\;\underbrace{\theta_{\text{global}}=\texttt{Agg}(\theta_{1},\theta_{2},\cdots,\theta_{C})}_{\text{Aggregate client model parameters on server}}\,,$
@@
where the client distributions may be homogeneous (i.e., $\forall\,(c,c^{\prime}),p_{c}=p_{c^{\prime}}$) or heterogeneous (i.e., $\exists\,c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$), Opt minimizes an objective $\mathcal{L}(\theta;p_{c})$ initialized at $\theta_{\text{init}}$. The most common objective in Federated learning is empirical risk minimization (ERM), which minimizing the average loss over the given dataset. Agg aggregates the client model parameters, where the most common aggregator is simply a (weighted) average of the client parameters, which corresponds to FedAvg (McMahan et al., 2017).
@@
Domain-based client heterogeneity: While previous client heterogeneity (i.e., $\exists\,c\neq c^{\prime},p_{c}\neq p_{c^{\prime}}$) is often expressed as label imbalance, i.e., $p_{c}(y)\neq p_{c^{\prime}}(y)$, we make a domain-based client heterogeneity assumption that each client distribution is a (different) mixture of train domain distributions, i.e., $p_{c}(x,y)=\sum_{d\in\mathcal{D}_{\text{train}}}w_{c,d}\,\,p_{d}(x,y)$ where $w_{c,d}$ is the weight of the $d$-th domain for the $c$-th client. At one extreme, FL with i.i.d. data would be equivalent to the mixture proportions being the same across all clients, i.e., $\forall c,c^{\prime},d,w_{c,d}=w_{c^{\prime},d}$, which we call the homogeneous setting. On the other extreme where the heterogeneity is maximum given the client number, we call complete heterogeneity (in Section 3. Table 1 summarizes the train-test heterogeneity in DG and the domain-based client heterogeneity from the FL context, where we focus on Federated DG.
@@
Table 1: Domain-based Federated DG considers the domain heterogeneity both among the clients’ local datasets (rows) and between the training and test datasets (columns). This paper focuses on the DG setting (right column).